TB, Inc. has a 20 percent tax rate and has $61.00 million in assets, currently financed entirely with equity. Equity is worth $5 per share, and book value of equity is equal to market value of equity. Also, let’s assume that the firm’s expected values for EBIT depend upon which state of the economy occurs this year, with the possible values of EBIT and their associated probabilities as shown below:

State Pessimistic Optimistic
Probability of state 0.45 0.55
Expected EBIT in state $3,644,750 $15,173,750

The firm is considering switching to a 30-percent-debt capital structure, and has determined that it would have to pay a 10 percent yield on perpetual debt in either event. What will be the level of expected EPS if GTB switches to the proposed capital structure? (Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

Expected EBIT = 0.45*3644750+0.55*15173750 = 9,985,700
Number of Shares before debt issue = Number of Shares/Price Per share = 61,000,000/5 = 12200000

Total Assets = 61 million
Debt amount after issue of debt =Total assets*30% = 61*30%=18.3 million
Interest Amount =Debt*Interest Rate =18.3*10% = 1.83 million
Amount of shares purchased = Debt Amount/Share price = 18,300,000/5 = 3660000

Expected EPS = (EBIT-Interest)*(1-Tax Rat)/(No of Shares before debt issue - Amount of shares purchased)
=(9985700-1830000)*(1-20%)/(12200000-3660000) = 0.764 or 0.76